OpenStudy (anonymous):

the midpoint of (2,3) and (5,7) is

hartnn (hartnn):

@claudineluya Hi :) \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

\[(\frac{ 5+2 }{ 2 },\frac{ 3+7 }{ 2 })\]\[(\frac{ 7 }{ 2 },5)\]

hartnn (hartnn):

@deena123 , you could have let @claudineluya do that work, after all its her/his question...

OpenStudy (anonymous):

..coordinating midpoint of (0,4) (0,8)

OpenStudy (anonymous):

\[(0,6)\]

OpenStudy (anonymous):

you can solve